40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 7 14:04:48 2011, steve, Update, SUS, He/Ne intensity noise effect on oplevs 7x
    Reply  Fri Oct 7 23:23:05 2011, kiwamu, Update, SUS, SRM oplev was oscillating 
       Reply  Wed Oct 19 10:49:35 2011, kiwamu, Update, SUS, SRM oscillation removed SRMoplevKWMtune.png
    Reply  Wed Oct 12 08:43:30 2011, steve, Update, SUS, ITMX oplev improved a bit ITMXoplev.pngITMXoplevservo_ON.png
       Reply  Thu Oct 13 14:23:09 2011, steve, Update, SUS, ITMX oplev with 3 mm beam on qpd ITMXoplev1103p#2.pngITMXoplev.png
          Reply  Mon Oct 17 17:07:30 2011, steve, Update, SUS, ETMX oplev returning beam od 3 mm ETMX-oplev.jpg
             Reply  Wed Oct 19 09:44:10 2011, steve, Update, SUS, ETMX oplev power spectrum ETMXoplev2lens.png
             Reply  Fri May 18 15:25:15 2012, steve, Update, SUS, AR coated lens swapped in at ETMX oplev ETMXoplevlensAR.png
Message ID: 6650     Entry time: Fri May 18 15:25:15 2012     In reply to: 5680
Author: steve 
Type: Update 
Category: SUS 
Subject: AR coated lens swapped in at ETMX oplev 

Quote:

 ETMX oplev had 6 mm diameter beam on the qpd.  I relayed the beam path with 2 lenses  to get back  3 mm beam on the qpd

BRC 037  -100 Bi _concave lens and PCX 25  200 VIS do the job. Unfortunately the concave lens has the AR 1064.

 

 

 The uncoated bi-concave lens was replaced by AR coated one: KBC 037 -100 AR.14 resulting 35% count increase on qpd

Attachment 1: ETMXoplevlensAR.png  1.103 MB  | Hide | Hide all
ETMXoplevlensAR.png
ELOG V3.1.3-